# The unified non-disciplinary model of the operation of neurons

### v3

@@ 5.2.1 Neuron as a PID controller
Biology uses a simple PID controller with proportional, integral, and derivative components. In simple words, it means that the output of the system (the output voltage measurable on the AIS) depends on the input voltage (the present: input current if the resistance is constant), the integrated input current(s) (the past), and their derivatives (the future). Notice that calculating the derivatives of the interdependent parameters needs care, as discussed in [62]. The sum of those summands defines the resulting output voltage. The biological case is more complicated than the technical one because biology works with slow ions, and the components have their temporal dependence (applying the laws created for fast currents needs emulation, as discussed in section 5), furthermore, the summands have several constituents and the neurons handle those current-related constituents autonomously. That is, unlike what is assumed in the classic physiology, the output voltage is not necessarily a simple function of the mentioned input variables; furthermore, it is not sufficient to consider the "present" of the system. The classic models all omit the derivatives’ contributions (a direct consequence of using ’clamping’: in that "frozen" state, the "future" cannot be interpreted) except that of the capacitive current, which is the passive consequence of the input currents. Furthermore, the ’Integrate and Fire’-type models consider the integrated currents only in the ’charge-up’ (’Computing’) period of operation, and they entirely omit that different physical processes are going on in different stages of operation. That is, even the best classic models can not describe neuronal operation completely and correctly.
@@
From a control theoretical point of view, the condenser geometry and the ion concentrations set the ’always the same’ value (the set-point; see also Figure 1) and the currents serve to keep or restore that value. After introducing a finite-width membrane into an electrolyte, a potential difference between the two membrane surfaces is created; see Eq. (19). When adjusting the membrane’s potential, we must consider its ground and excited states separately, given that the perturbations in these two cases are vastly different. Nature employs various mechanisms to maintain the ground state and recover it after generating a spike. Although nature’s tools are remarkably similar (channels and pumps; here, for discussing ’downhill’ channels, we must consider the charge-up of different components), the significant difference in current intensity warrants a separate discussion.
@@ 6.1 The physical process
In the transient (perturbed) state, the process variable (the membrane’s potential) exceeds the threshold. That event acts as a fast trigger signal. The amount of charged particles (and so: the membrane’s voltage) suddenly increases, and the created gradient drives a current. When $Na^{+}$ ions rush into the intracellular layer, they roughly increase the overall concentration and the potential in that thin layer. All other ions, including $K^{+}$, also feel a driving force. The targeted membrane potential is set electrically, and the driving gradients may behave in unexpected ways during the transient period. The actual voltage gradient may temporarily reverse the direction of the chemical gradients.
@@
Figure 12: A summary of producing neuronal AP as a 3-D figure. The thick, solid 3-D line shows how the color-coded voltage gradient drives the membrane potential over time. The three dashed diagram lines are the projections of the spatial curve to the boundary planes of the cube. The physiology delivers 2-D plots; the figure explains how those parameters are interrelated. The red line shows the voltage gradient produced by the rush-in of the $Na^{+}$ ions, see also the middle inset in Fig. 13. It produces the blue diagram line AP, see also the bottom inset in Fig. 13. The interdependence of the gradient and the voltage it produces leads to Fig. 15, left side, that allows to calculate the thermodynamic characteristics, right side, of neuronal operation.
@@ 6.4 Damped oscillation
The "moving force" has two major components. The sudden rush-in of $Na^{+}$ ions has (at least) a double effect. One effect is that they press the surface of the membrane that elastically increases its diameter [80]. Due to their mutual repulsion, the ions experience a strong electrical force toward the membrane. It represents an impulse $J=F\,\Delta t$ (N  s), that enables estimating the time and energy of the change the rush-in causes and explains that the energy needed to generate an AP is suddenly produced electrically, and stored temporarily as elastic energy (the neuron produces the required energy in the background, when ATP produces ions by hydrolysis, as described above); see Appendix B. The case is practically identical to the deformation of an elastic plate induced by the hydrostatic pressure of a water column. At the beginning, the aluminum plate is suddenly subjected to hydrostatic pressure, and it reaches equilibrium after initial oscillations. The diagram line of the process is shown in the figure [81], and the method of simulation is discussed in [82]. Compare the diagram line to the gradient in Fig. 13, middle inset.
@@ 7.1 Deriving thermodynamic description
As the middle inset in Fig. 13 shows, the sum of the rush-in and AIS gradients (plus more gradients if any) controls the generation of the AP. During excitation, $Q_{X}$ energy is fed into the neuron. The rush-in of the $Na^{+}$ ions injects more $Q_{R}$ energy, which can be used to generate an AP. It is a kind of potential energy, which is converted in the first phase of $AP$ (while the voltage gradient is positive) to the kinetic energy $Q_{H}$ of AP. When the membrane potential reaches its peak, the gradient reverses, decreasing the potential. In the second phase of AP (while the voltage gradient is negative) the neuron harvests $Q_{L}$ energy. At the beginning of the transient state, the size of the neuron (due to the increased electrostatic repulsion) slightly increases [80] and stores the energy mainly in the form of elastic energy. The resultant of the electrical, elastic, and thermodynamic forces moves the ions toward the AIS. Although part of the energy is dissipated while the ions exit through the AIS with their Stokes-Einstein speed, the rest of the energy (the elastic part) is recovered. The concentration and potential values change as discussed in connection with Fig. 10. The membrane voltage controls synapses as shown in the top inset in Fig. 13.
@@
Figure 16: The time course of the electrical gradient [53] perfectly matches that of the mechanical and thermal changes [35]; however, classic theory cannot explain the connection between them [31]. The experimental $\frac{dV}{dt}$ uses an opposite sign convention.
@@
In terms of thermodynamics, the ions stop in the immediate vicinity of the membrane, i.e., initially, the pressure created by their mutual repulsion is present at the inner surface of the sphere (it is a mechanical shock wave). In the positive phase of the damped oscillation, it compresses the fluid in the sphere. In the second phase, in the negative phase of oscillation, it decompresses the fluid (sucks back the fluid). Given that the density of ions in the volume is constant (the ions’ speed is too low to follow those sudden changes), the electrical and thermodynamic quantities faithfully follow each other. Measuring electrical potential and mechanical pressure yields disciplinary measurements of the same physical effect.
@@
From Fig. 15, after recalibration (with $R_{AIS}=10^{7}$ Ohm), one can estimate (by approximating roughly the area under the arcs by triangles, in units of $10^{-7}[J]$), $Q_{H}=5.2,Q_{L}=1.8,W=3.4,Q_{X}=0.2$. That is, according to the "phase plot" [53], the neuron operates with an efficiency around 65%, and approximately 3% of the energy is derived from the upstream neurons. The values (without using data from dedicated measurements) agree excellently with the measured [43] values $W=2.5$ and 74%; furthermore, the $Q_{X}$:total ratio $1:35$, derived by measuring the direct energy consumption [45]. An interesting proof that the two disciplines see the same physical effect, that electrical parameters can measure the thermodynamic efficiency. The invested energy can be safely estimated from the work the rush-in charge performs given that the energy is almost entirely stored as elastic potential energy due to the enormous elasticity modulus
7.2 Heat emission/absorption
In the framework of the classic theory, Ohmic currents flow through resistors that dissipate heat due to friction, no matter in which direction ($W=I^{2}\times R$) the ion currents flow. The existence of "cooling", alone, undermines the credibility of the classic theory [66]. The issue here is that the dissipation was attributed to the "leakage current." As Hodgkin wrote: "Hill and his colleagues found [22] that an initial phase of heat liberation [of the action potential] was followed by one of heat absorption. […] a net cooling on open-circuit was totally unexpected and has so far received no satisfactory explanation." [83] "All authors came to similar conclusions: during the action potential, no significant net heat is produced. Transient heat releases are mostly reabsorbed in the second phase of the action potential. …The finding of a reversible heat release during the action potential of nerves is a striking and very fundamental fact. It is inconsistent with the HH model. The physics underlying the nervous impulse must rather be based on reversible processes". [84]
Heat liberation and absorption are direct experimental evidence of the "slow current" and the mechanisms described above. The lack of leakage current (see section 4.5.4) not only validates our model, but also solves the long-standing mystery of reversible heat release during the AP. Pressure changes accompany the electrical process and naturally account for the observed temperature changes. Our Speculations in Appendix B on the energy consumption of the nervous impulse underpins that the complex physical process comprises mostly reversible disciplinary processes: most of the energy of AP is stored as reversible elastic potential energy. Hence, its dissipation is also negligible: "no significant net heat is produced" [84]. (Alternatively, it may be hypothesized that the positive $K^{+}$ ions, while traveling ahead of the positive $Na^{+}$ ions under the effect of a potential-free space in the second phase of AP, not only turn with their positive charge the offset voltage to negative, but on the way the $K^{+}$ ions collect the heat dissipated by the $Na^{+}$ ions. "Laws of science cannot describe life".)
The shocking conclusion in [63], that "brain cells communicate with mechanical pulses, not electrical signals", more precisely sounds that mechanical pulses (pressure waves) deliver the charge carriers of the electrical signals, in close interaction with the repulsion of the delivered charged carriers. In Equation (7), we also must consider the mechanical force delivered by the pressure wave. Furthermore, the nervous pulse transmission over the axons must be revisited. The classical mechanism based on coordinated action of ion channels is undoubtedly wrong. The enormous pressure, the incompressible electrical fluid, and the repulsion of the delivered particles suggest an alternative and more reasonable mechanism.
@@ 8 Summary
The paper reviewed the competing disciplinary electrical and thermodynamic theories for describing neuronal operation and described the non-disciplinary unified theory that resolves all former controversies between the two. We have shown that the recent incompatibility between the electrical and thermodynamic theories is the restrictive disciplinary approach on both sides, and that using a cross-disciplinary approach (and revisiting the fundamental laws of science for describing living matter) perfectly describe neuronal operation. We have shown that the interplay of those two disciplines is vital in all aspects of neuronal operation. The neuron represents a well-designed, simple control circuit with separate ground and excited state mechanisms, where the construction defines the setpoint and the control is implemented by regulating the ion concentrations on the two sides of the semipermeable membrane. The geometry and the overall concentration of the cellular fluids define the nominal values of the electrical and thermal gradients. The AP is part of the operation of the control circuit. The low and largely variable speed of ionic current, in which the discrete features of the carrier dominate, creates local ionic gradients and movement in the electrolyte in the segments of the neuron.
@@
Slow currents deliver ions through the permanently open "resting" and the gated "transient" ion channels, respectively, which means they affect the gradients until the equilibrium state is kept or restored. By introducing the theoretical idea that biological currents are slow, we underpinned the experimental observation by Hodgkin and Huxley [74] that "the changes in ionic permeability depend on the movement of some component of the membrane which behaves as though it had a large charge or dipole moment". By using the principles of classical electricity, we calculated, that neurons handle charge as electricity describes (electrolyte) condensers. We defied the hypothesis suggested by HH that some "leakage current" produces the resting potential; this way solved the decades-old mystery of "heat production" by neurons and the controversy of the experienced and theoretically expected energy consumption of neural computing. We fixed the mistakes HH claimed in the physical picture behind their empirical description and mathematical equations, and introduced the correct physical model by providing its detailed mathematical description. By understanding that the biological "construction" defines the value of the resting potential, we defied the decades-old fallacy that a mystic concerted operation of ion channels and a linear combination of poorly-defined permeability values (the GHK equation) define the resting potential.
@@
We could theoretically underpin (see Fig. 9) that "action potential generation in nearly all types of neurons and muscle cells is accomplished through mechanisms similar to those first detailed in the squid giant axon in early research by HH[17]. However, not because "this unequal distribution of ions is maintained by ionic pumps and exchangers". Instead, electrostatic charging is due to charge separation at the membrane and to polarization in the electrolyte surrounding it. The mechanically separated segments produce two charged layers in the electrolyte proximal to the membrane. Those layers have a decisive impact on the neuron’s electrical operation.
@@
The operation can be described in terms of electricity and thermodynamics, although the ’non-ordinary’ laws [31] of science must be considered. We successfully solved the mystery of the heat absorption of the neuron and derived a method for the thermodynamic description of the neuron (albeit based on parameters that can be measured electrically). Furthermore, its energy consumption and operational efficiency. We showed that, due to the disciplinary properties, one cannot derive the thermodynamic entropy using the usual methods of information processing.
@@
The research did nor receive any support.
@@ A.1 Science methods
For centuries, science developed its methods for deriving abstract concepts by reducing the features of a real object to an abstract one that cannot be reduced further, such as mass and charge. It derived laws for the forces acting on those abstract objects, such as Newton’s universal law of gravitation and Coulomb’s law of electricity. Then one could apply Newton’s laws of motion. Experience shows that the generated forces, independently of their origin, can be summed, and one can apply the laws of motion by using the resultant force. Ions are special from the point of view of reducing the ’material point’ to one single abstraction: the ions are charge and mass simultaneously, without a further possibility of reduction. Notice that in the electrical abstraction, no mass is present, so one can use the equations assuming ’instant interaction’, which in biology lead to non-physical explanations of the observations (such as ’delayed current’). In the mechanical/thermodynamic abstraction, mass is to be moved, making the finite-speed interaction evident. The same physical phenomenon, the interaction (or movement) of ions is described using an ’infinite’ electrical speed and a million times less mass propagation speed, respectively, which leads to an unresolvable discrepancy, given that physics is not prepared to handle different speeds in the same interaction event [6]. This item is listed in Table 1 in [9] as the "Electrical vs Mechanical" dichotomy.
The low speed of ions in electrolytes introduces further problems. The farther parts of the cell will see any change in the local value of the state variables with a delay. The material transport represents simultaneously mass and charge, so the transport itself changes the gradients. This process keeps the entire volume of the electrolyte in (more or less) continuous change (which makes life). Furthermore, biological objects inside the cell can absorb ions and charge up. With their potential, they alter local gradients, accelerating or decelerating ions. Not to mention that biological objects can be active in the sense that (depending on the environmental conditions) they can let ions from one separated volume part into the other.
@@
Although science is aware that the apparently continuous matter in nature cannot be divided infinitely, even though it knows that the abstracted discrete ’material points’ (as A. Einstein coined) have well-defined discrete values, it occurs rarely that both views must be applied in describing a single phenomenon. The continuous and discrete approaches (also called macroscopic and microscopic views) seem to be independent of each other. Connecting those views was one of the tasks performed by thermodynamics for particles with no long-range interactions. However, there is no similar discipline for electricity, although the behavior of charge carriers is similar to that of neutral discrete particles. It is not evident for electrons, but significant for ions. This item is listed in Table 1 in [9] as the "Macroscopic vs Microscopic" dichotomy.
@@
The dichotomy about reversibility is closely related to the exclusively electrical nature of the HH theory. In the framework of that theory, Ohmic currents flow through resistors that irreversibly dissipate heat due to friction, no matter in which direction ($W=I^{2}\times R$) the ion currents flow. By introducing the concept of ’delayed current’ and the hypothesis that some hidden power controls the operation of neurons by changing their conductance, physiology gave rise to the fallacy that science and life sciences are almost exclusive fields. Furthermore, their (unintended) model provokes questions (for a review see [12]) whether it is a model at all, and what controversies it delivers. As Hodgkin wrote [83]: "Hill and his colleagues found [22] that an initial phase [of the action potential] was followed by one of heat absorption. […] a net cooling on open-circuit was totally unexpected and has so far received no satisfactory explanation." Since that invention, experimental evidence from the missing "leakage current" [45] to the conversion between elastic and kinetic energies [65] to demonstrating the pressure-wave-like behavior of AP witnesses that at least part of the phenomenon is of thermodynamic origin; that is, it requires a cross-disciplinary explanation. For discovering the reciprocal relations in thermodynamics [87], also in electrolytes, Lars Onsager was awarded the 1968 Nobel Prize in Chemistry. Those relations, together with the theoretical understanding, manifesting in the cross-disciplinary Nernst-Planck relation, paved the way to a combined theory. However, no thermodynamic explanation has been given so far. Our discussion provides, in line with the experimental results, a complex electrical/thermodynamic description of neuronal operation.
@@
Physics notoriously suffers from a lack of handling different simultaneous interactions; facing such a case leads to misunderstandings, debates, and causality problems. Such a famous case is the entanglement speed. In that time, E. Schrödinger introduced his famous law of motion in quantum mechanics entirely analogously to how I. Newton introduced his Laws of Motion. Similar to the Newtonian ’absolute time’, the quantum mechanical interaction is supposed to be ’instant’ (this is the price for having ’nice’ equations in classical and quantum mechanics), i.e., its speed is supposed to be infinitely high. However, at that time, it was already known that the speed of electrical interaction (propagation of electromagnetic waves) is finite. So if an object has quantum-mechanical interaction (aka entanglement) and electrical interaction simultaneously, the corresponding forces act simultaneously but reach the other object at different times. The entanglement arrives instantly; the electromagnetic effect arrives at a time we can calculate from the interaction speed and the spatial distance between the objects. This effect leads to causality problems: the two interactions of photons entangled earlier in an exploded supernova should be measured at two different times, meaning a "spooky remote interaction" as A. Einstein coined, and leads to contradictions such as the Einstein–Podolsky–Rosen paradox. Actually, the issue stems from the improper handling of mixing interaction speeds: the Schrödinger-equation introduces the infinitely large interaction speed, while the electromagnetic (EM) interaction has a finite speed.

### v4

@@ 5.2.1 Neuron as a PID controller
Biology uses a simple PID controller with proportional, integral, and derivative components. In simple words, it means that the output of the system (the output voltage measurable on the AIS) depends on the input voltage (the present: input current if the resistance is constant), the integrated input current(s) (the past), and their derivatives (the future). Notice that calculating the derivatives of the interdependent parameters needs care, as discussed in [29]. The sum of those summands defines the resulting output voltage. The biological case is more complicated than the technical one because biology works with slow ions, and the components have their temporal dependence (applying the laws created for fast currents requires emulation, as discussed in section 5). Furthermore, the summands have several constituents, and the neurons handle those current-related constituents autonomously. That is, unlike what is assumed in the classic physiology, the output voltage is not necessarily a simple function of the mentioned input variables. Furthermore, it is not sufficient to consider the "present" of the system. The classic models all omit the derivatives’ contributions (a direct consequence of using ’clamping’: in that "frozen" state, the "future" cannot be interpreted) except that of the capacitive current, which is the passive consequence of the input currents. Furthermore, the ’Integrate and Fire’-type models consider the integrated currents only in the ’charge-up’ (’Computing’) period of operation, and they entirely omit that different physical processes are going on in different stages of operation. That is, even the best classic models can not describe neuronal operation completely and correctly.
@@
From a control theoretical point of view, the condenser geometry and the ion concentrations set the ’always the same’ value (the setpoint; see also Figure 2) and the currents serve to keep or restore that value. After introducing a finite-width membrane into an electrolyte, a potential difference between the two membrane surfaces is created; see Eq. (21). When adjusting the membrane’s potential, we must consider its ground and excited states separately, given that the perturbations in these two cases are vastly different. Nature employs various mechanisms to maintain the ground state and recover it after generating a spike. Although nature’s tools are remarkably similar (channels and pumps; here, for discussing ’downhill’ channels, we must consider the charge-up of different components), the significant difference in current intensity warrants a separate discussion.
@@ 6.1 The physical process
In the transient (perturbed) state, the process variable (the membrane’s potential) exceeds the threshold. That event acts as a fast trigger signal. The amount of charged particles (and so: the membrane’s voltage) suddenly increases, the created gradient drives a current. When $Na^{+}$ ions rush into the intracellular layer, they roughly increase the overall concentration and the potential in that thin layer. All other ions, including $K^{+}$, also feel a driving force. The targeted membrane potential is set electrically, and the driving gradients may behave unexpectedly during the transient period. The actual voltage gradient may temporarily reverse the direction of the chemical gradients.
@@
Figure 13: A summary of producing neuronal AP as a 3-D figure. The thick, solid 3-D line shows how the color-coded voltage gradient drives the membrane potential over time. The three dashed diagram lines are the projections of the spatial curve to the boundary planes of the cube. The experimental physiology delivers those 2-D plots; the figure explains how those parameters are interrelated. The red line shows the voltage gradient produced by the rush-in of the $Na^{+}$ ions, see also the middle inset in Fig. 14. It produces the blue diagram line AP, see also the bottom inset in Fig. 14. The interdependence of the gradient and the voltage it produces leads to Fig. 17, left side, that allows to calculate the thermodynamic characteristics, right side, of neuronal operation.
@@ 6.3 Mathematics of the ActionPotential
Fig. 15 compares the calculated function shape calculated by Eq. (50) to the function fitted by HH to their experimental observations.
Figure 15: Comparing the calculated AP to the empirical function fitted by Hodgkin-Huxley (Fig. 17 of   [16]).
@@ 6.4 Damped oscillation
The "moving force" has two major components. The sudden rush-in of $Na^{+}$ ions has (at least) a double effect. One effect is that they press the surface of the membrane that elastically increases its diameter [31]. Due to their mutual repulsion, the ions experience a strong electrical force toward the membrane. It represents an impulse $J=F\,\Delta t$ (N  s), that enables estimating the time and energy of the change the rush-in causes and explains that the energy needed to generate an AP is suddenly produced electrically, and stored temporarily as elastic energy (the neuron produces the required energy in the background, when ATP produces ions by hydrolysis, as described above). The case is practically identical to the deformation of an elastic plate induced by the hydrostatic pressure of a water column. At the beginning, the aluminum plate is suddenly subjected to hydrostatic pressure, and it reaches equilibrium after initial oscillations. The diagram line of the process is shown in the figure [90], and the method of simulation is discussed in [91]. Compare the diagram line to the gradient in Fig. 14, middle inset.
@@ 7 Thermodynamics of Action Potential
As discussed above, by finding the connection between apparently separated disciplines, one can derive from quantities belonging to the discipline electricity quantities belonging to thermodynamics; see section 3.4. In this section, we derive the description of the thermodynamic Carnot cycle from measured electrophysical data, essentially from the 2-D diagram lines shown in Fig. 13.
@@ 7.1 Deriving thermodynamic description
As the middle inset in Fig. 14 shows, the sum of the rush-in and AIS gradients (plus more gradients if any) controls the generation of the AP. During excitation, see Fig. 17, $Q_{X}$ energy is fed into the neuron. The rush-in of the $Na^{+}$ ions injects more $Q_{R}$ energy, which can be used to generate an AP. It is a kind of potential energy, which is converted in the first phase of $AP$ (while the voltage gradient is positive) to the kinetic energy $Q_{H}$ of AP. When the membrane potential reaches its peak, the gradient reverses, decreasing the potential. In the second phase of AP (while the voltage gradient is negative) the neuron harvests $Q_{L}$ energy. At the beginning of the transient state, the size of the neuron (due to the increased electrostatic repulsion) slightly increases [31] and stores the energy mainly in the form of elastic energy. The resultant of the electrical, elastic, and thermodynamic forces moves the ions toward the AIS; see Fig. 1. Although part of the energy is dissipated while the ions exit through the AIS with their Stokes-Einstein speed, the rest of the energy (the elastic part) is recovered. The concentration and potential values change as discussed in connection with Fig. 11. The membrane voltage controls synapses as shown in the top inset in Fig. 14.
@@
Figure 18: The red diagram lines show the membrane voltage and its gradient during producing an action potential [67], Fig. 4. The black diagram lines show the size-related and heat-related changes during producing an action potential [52], Fig. 1. and [40], Fig. 18.5.
The time course of the electrical gradient [67] perfectly matches that of the mechanical and thermal changes [52]; however, classic theory cannot explain the connection between them [48]. The experimental $\frac{dV}{dt}$ uses an opposite sign convention.
@@
In terms of thermodynamics, the ions stop in the immediate vicinity of the membrane, i.e., initially, the pressure created by their mutual repulsion is present at the inner surface of the sphere (it is a mechanical shock wave). In the positive phase of the damped oscillation, it compresses the fluid in the sphere. In the second phase, in the negative phase of oscillation, it decompresses the fluid (sucks back the fluid). Given that the density of ions in the volume is constant (the ions’ speed is too low to follow those sudden changes), the electrical and thermodynamic quantities faithfully follow each other. Measuring electrical potential and mechanical pressure yields measurements of the same physical effect within the same discipline.
@@
From Fig. 17, after recalibration (with $R_{AIS}=10^{7}$ Ohm), one can estimate (by approximating roughly the area under the arcs by triangles, in units of $10^{-7}[J]$), $Q_{H}=5.2,Q_{L}=1.8,W=3.4,Q_{X}=0.2$. That is, according to the "phase plot" [67], the neuron operates with an efficiency around 65%, and approximately 3% of the energy is derived from the upstream neurons. The values (without using data from dedicated measurements and making precise numerical integration) agree excellently with the measured [36] values $W=2.5$ and 74%; furthermore, the $Q_{X}$:total ratio $1:35$, derived by measuring the direct energy consumption [35]. An interesting proof that the two disciplines see the same physical effect, that electrical parameters can measure the thermodynamic efficiency. The invested energy can be safely estimated from the work the rush-in charge performs, given that the energy is almost entirely stored as elastic potential energy due to the enormous elasticity modulus
@@ 8 Summary
The paper reviewed the competing disciplinary electrical and thermodynamic theories for describing neuronal operation and describing their synthesis, the non-disciplinary unified theory that resolves all former controversies between the two. The interplay of those two disciplines is vital in all aspects of neuronal operation. The restrictive disciplinary approach on both sides causes the apparent incompatibility between them. Using a cross-disciplinary approach (and revisiting the fundamental laws of science for describing living matter) one can perfectly describe neuronal operation. The neuron represents a well-designed, simple control circuit with separate ground and excited state mechanisms, where the construction defines the setpoint and the control is implemented by regulating the ion concentrations on the two sides of the semipermeable membrane. The geometry and the overall concentrations of the cellular fluids define the nominal values of the electrical and thermal gradients. The AP is part of the operation of the control circuit. The low and largely variable speed of ionic currents, in which the discrete features of the carrier dominate, creates local ionic gradients and movement in the electrolyte in the segments of the neuron.
@@
Slow currents deliver ions through the permanently open "resting" and the gated "transient" ion channels, respectively, which means they affect the gradients until the equilibrium state is kept or restored. By introducing the theoretical idea that biological currents are slow, we underpinned the experimental observation by Hodgkin and Huxley [84] that "the changes in ionic permeability depend on the movement of some component of the membrane which behaves as though it had a large charge or dipole moment". Using the principles of classical electricity, we calculated that neurons handle charge as electricity describes (electrolyte) condensers. Given the presence of electrolytes on both sides of the membrane, the correct values for the resting potential and the time course of the action potential have been derived numerically from first principles, without empirical functions. We defied the hypothesis suggested by HH that some "leakage current" produces the resting potential; this way solved the decades-old mystery of "heat production" by neurons and the controversy of the experienced and theoretically expected energy consumption of neural computing. We fixed the mistakes HH claimed in the physical picture behind their empirical description and mathematical equations, and introduced the correct physical model by providing its detailed mathematical description. By understanding that the biological "construction" defines the value of the resting potential, we defied the decades-old fallacy that a mystic concerted operation of ion channels and a linear combination of poorly-defined permeability values (the GHK equation) define the resting potential.
@@
We could theoretically underpin that "action potential generation in nearly all types of neurons and muscle cells is accomplished through mechanisms similar to those first detailed in the squid giant axon in early research by HH[19]. However, not because "this unequal distribution of ions is maintained by ionic pumps and exchangers". Instead, electrostatic charging is due to charge separation at the membrane and to polarization in the electrolyte surrounding it. The mechanically separated segments produce two charged layers in the electrolyte proximal to the membrane. Those layers have a decisive impact on the neuron’s electrical operation.
@@
The operation can be described in terms of electricity and thermodynamics, although the ’non-ordinary’ laws [48] of science must be considered. We solved the mystery of the neuron’s heat absorption and derived a method for the thermodynamic description of the neuron, including its energy consumption and operational efficiency. We demonstrated the need for a cross-disciplinary approach by calculating the thermodynamic parameters from experimental data that can be measured electrically. We showed that, due to the disciplinary properties, one cannot derive the thermodynamic entropy using the usual methods of information processing.
@@
The research did not receive any support.
@@ A.1 Science methods
For centuries, science has developed its methods for deriving abstract concepts by reducing the features of a real object to an abstract one that cannot be reduced further, such as mass and charge. It derived laws for the forces acting on those abstract objects, such as Newton’s universal law of gravitation and Coulomb’s law of electricity. Then one could apply Newton’s laws of motion. Experience shows that the generated forces, independently of their origin, can be summed, and one can apply the laws of motion by using the resultant force. Ions are special from the point of view of reducing the ’material point’ to one single abstraction: the ions are charge and mass simultaneously, without a further possibility of reduction. Notice that in the electrical abstraction, no mass is present, so one can use the equations assuming ’instant interaction’, which in biology led to non-physical explanations of the observations (such as ’delayed current’). In the mechanical/thermodynamic abstraction, mass is to be moved, making the finite-speed interaction evident. The same physical phenomenon, the interaction (or movement) of ions is described using an ’infinite’ electrical speed and a million times less mass propagation speed, respectively, which leads to an unresolvable discrepancy, given that physics is not prepared to handle different speeds in the same interaction event [8]. This item is listed in Table 1 in [11] as the "Electrical vs Mechanical" dichotomy.
The low speed of ions in electrolytes introduces further problems. The farther parts of the cell will see any change in the local value of the state variables with a delay. The material transport represents simultaneously mass and charge, so the transport itself changes the gradients. This process keeps the entire volume of the electrolyte in (more or less) continuous change (that makes life). Furthermore, biological objects inside the cell can absorb ions and charge up. With their potential, they alter local gradients, accelerating or decelerating ions. Not to mention that biological objects can be active in the sense that (depending on the environmental conditions) they can let ions from one separated volume part into the other.
@@
Although science is aware that the apparently continuous matter in nature cannot be divided infinitely, even though it knows that the abstracted discrete ’material points’ (as A. Einstein coined) have well-defined discrete values, it occurs rarely that both views must be applied in describing a single phenomenon. The continuous and discrete approaches (also called macroscopic and microscopic views) seem to be independent of each other. Connecting those views was one of the tasks performed by thermodynamics for particles with no long-range interactions. However, there is no similar discipline for electricity, although the behavior of charge carriers is similar to that of neutral discrete particles. That effect is not evident for electrons, but significant for ions. This item is listed in Table 1 in [11] as the "Macroscopic vs Microscopic" dichotomy.
@@
The dichotomy about reversibility is closely related to the exclusively electrical nature of the HH theory. In the framework of that theory, Ohmic currents flow through resistors that irreversibly dissipate heat due to friction, no matter in which direction ($W=I^{2}\times R$) the ion currents flow. By introducing the concept of ’delayed current’ and the hypothesis that some hidden power controls the operation of neurons by altering their conductance, physiology gave rise to the fallacy that the sciences and life sciences are almost exclusive fields. Furthermore, their (unintended) model provokes questions (for a review see [14]) whether it is a model at all, and what controversies it delivers. As Hodgkin wrote [39]: "Hill and his colleagues found [26] that an initial phase [of the action potential] was followed by one of heat absorption. […] a net cooling on open-circuit was totally unexpected and has so far received no satisfactory explanation." Since that invention, experimental evidence from the missing "leakage current" [35] to the conversion between elastic and kinetic energies [32] to demonstrating the pressure-wave-like behavior of AP witnesses that at least part of the phenomenon is of thermodynamic origin; that is, it requires a cross-disciplinary explanation. For discovering the reciprocal relations in thermodynamics [94], also in electrolytes, Lars Onsager was awarded the 1968 Nobel Prize in Chemistry. Those relations, together with the theoretical understanding, manifesting in the cross-disciplinary Nernst-Planck relation, paved the way to a combined theory. However, no thermodynamic explanation has been given so far. Our discussion provides, in line with the experimental results, a complex electrical/thermodynamic description of neuronal operation.
@@
Physics notoriously suffers from a lack of handling different simultaneous interactions; facing such a case leads to misunderstandings, debates, and causality problems. Such a famous case is the entanglement speed. In that time, E. Schrödinger introduced his famous law of motion in quantum mechanics entirely analogously to how I. Newton introduced his Laws of Motion. Similar to the Newtonian ’absolute time’, the quantum mechanical interaction is supposed to be ’instant’ (this is the price for having ’nice’ equations in classical and quantum mechanics), i.e., its speed is supposed to be infinitely high. However, at that time, it was already known that the speed of electrical interaction (propagation of electromagnetic waves) is finite. So if an object has quantum-mechanical interaction (aka entanglement) and electrical interaction simultaneously, the corresponding forces act simultaneously but reach the other object at different times. The entanglement arrives instantly; the electromagnetic effect arrives at a time we can calculate from the interaction speed and the spatial distance between the objects. This effect leads to causality problems: the two interactions of photons entangled earlier in an exploded supernova should be measured at two different times, meaning a "spooky remote interaction" as A. Einstein coined, and leads to contradictions such as the Einstein–Podolsky–Rosen paradox. Actually, the issue stems from the improper handling of mixing interaction speeds: the Schrödinger-equation introduces the infinitely large interaction speed, while the EM interaction has a finite speed.